# Dynamical and structural properties of an absorbing phase transition: a case study from granular systems

### v1

@@ IV.1 Simple $\Delta$ model
Figure 8: Comparison between the different theories applied to the simple $\Delta$ model without synchronization and simulations. The analysis is accompanied by examples of the breakdown of the assumptions used. a) Comparison of the theory of the simple model with the simulations. “Naive theo.” is Eq. (15). “Corrected theo.” is Eqs. (18) and (20). Dark dashed lines are the theoretical critical packing fractions derived from the mean free path argument. The inset is in a semi-log scale close to the transition and in units of $m\Delta^{2}/2$. We see that when the naive theory predicts a temperature of the order $m\Delta^{2}/2$, the system reaches an absorbing state in simulation. $N=5000$, $\alpha=0.95$ and $\hat{\tau}\gamma=0.01$. The increasing values of $\hat{\tau}\Delta/\sigma$ are: 0.0125, 0.015 and 0.0175. b) Evolution of the velocity autocorrelation function as the critical density of the continuous transition is approached. The system starts to exhibit backscattering and the prediction from the Enskog theory with molecular chaos cannot be trusted. c) Evolution of the pre-colliding velocity $v_{coll}$ as a function of the proximity to the absorbing state. $v_{coll}$ tends to saturate, implying that the active particles have a finite energy at the transition. d) Evolution of the measured frequency of collision as the critical point is approached ($T^{ss}\rightarrow 0$) by varying $\Delta$ from $\Delta_{c}$, which gives a vanishing $T^{ss}$ to a large value where the measured frequency of collision matches the one predicted by the kinetic theory $\omega_{e}$. At small temperatures, the frequency of collision increases linearly with $T^{ss}$ because it becomes proportional to the number of active particles.
Before introducing the theory of the full coarse-grained model, it is useful to discuss the scenario where particles never synchronize which corresponds to the case $\tau_{s}\rightarrow\infty$ and, consequently, $\Delta_{ij}=\Delta$ is constant, see Eq (7). If the theory is accurate it should predict, in this situation, a continuous transition. We first recall the theory developed in Ref. 53 and introduced in Ref. 92.
@@
Using the collision rule of the simple $\Delta$ model (Eq. (5)) and averaging over collisions, the energy change per collision can be estimated explicitly [92]:
@@
where $m$ is the mass of a particle. This result is based on the assumption of uncorrelated velocities (molecular chaos) and Gaussianity of the velocity distribution. The detailed calculations required to arrive at Eq. (12) can be found in Ref. 53. These assumptions immediately imply that the frequency of collision is given by the so-called Enskog frequency of collision $\omega_{e}$ [126]:
@@
where $l(\phi)$ is the mean free path and $\chi$ is the Enskog factor taken to be the radial pair distribution function at contact of an equilibrium system at the same density. It arises from the Enskog’s correction to the molecular chaos assumption [126]. It can be approximated for $\phi<0.6$ by [127]:
@@
In Fig. 8a, we compare this theoretical prediction with measurements done with simulations of the coarse-grained model at fixed $\Delta$ (i.e. no synchronization) and indeed the theory seems to work very well. This result, however, must be taken with care since the simulations reach an absorbing state at a finite packing fraction and this behavior is not predicted by the theory, which predicts zero temperature only at zero packing fraction, see the inset of Fig. 8a. The assumptions of homogeneity, molecular chaos and Gaussian distribution of velocity do not hold close to or below the critical packing fraction due to the significant effect of viscous drag. As can be seen in Fig. 8b, close to the transition, molecular chaos is broken since the velocity autocorrelation function develops a negative minimum at a time roughly equal to the mean collision time that is a clear indication of a backscattering effect. This is incompatible with the assumption of uncorrelated or Markovian collisions underlying the molecular chaos assumption. Indeed, such a theory can only predict exponentially decaying velocity autocorrelation [128, 129, 130]. This backscattering is induced by the considerable slowdown of the particles before colliding with their neighbors and gaining velocities in a direction opposite to the one they came from. Treating this effect would require going beyond molecular chaos; an approach that is made possible, by some tedious calculation [131]. Moreover, we observed in Fig. 7 that close to the critical packing fraction, the dynamics is dominated by avalanches of active particles within a population of inactive particles. The transition could then occur via two different scenarios. Either the temperature of the active particles reaches 0 at $\phi_{c}$, or their number vanishes while they still possess a finite amount of energy. The second scenario is the one observed in our system, as illustrated in Fig. 8c. Indeed, as the critical point is approached, the pre-colliding velocity $v_{0}$ (the velocity just before a collision) tends to saturate. This indicated that, at the transition, the active particles have finite energy, roughly equal to the minimal excitation possible: $m\Delta^{2}/2$, but their number vanishes.
@@
The first term accounts for the conversion of inactive particles to active ones due to collisions. The second term is a rate of inactivation modulated by the probability for an active particle of not colliding in a time $\tau_{a}$ assuming that collisions are Poissonian [132]. It accounts for particles reaching the absorbing state due to an absence of collision between them and a neighbor in a time $\tau_{a}$. This $\tau_{a}$ is acting as a cutoff with a value that will be derived below and should be inversely proportional to the drag $\gamma$. We recall that $\omega/2$ is the number of collision events per particle happening in the system per unit of time, while $\omega$ is the average frequency at which one particle collides with others.
@@
with $W_{0}$ the principal branch of the Lambert $W$ function. $P_{a}^{ss}$ varies continuously, and the transition happens when the mean free time $1/\omega_{e}$ of the active particle is of the same order ($W_{0}(2)\simeq 0.85)$ as the stopping time $\tau_{a}$. It is gratifying to observe that Eq. (19) is the mean field equation describing the conserved directed percolation universality class, which is the correct equation to describe the vicinity of our phase transition [114]. At this point, it is interesting to make a parallel with models on lattice. As said before, the criticality in our model comes from the vanishing of $P_{a}$ while the temperature of the active particles is fixed. This behavior closely resembles lattice models in the conserved directed percolation universality class, where the order parameter is the number of active particles and for which the activity kick allowing a particle to move on the lattice is fixed and finite [107, 133].
To close our equations, we must determine a suitable value for $\tau_{a}$. This can be achieved, for example, by finding an independent prediction for $\phi_{c}$. An argument presented in Ref. [93] allows us to find a lower bound for the critical packing fraction of the continuous transition. If we assume spatial homogeneity, at the critical packing fraction, particles must collide on average with zero velocity. Below this density, particles do not have sufficient momentum after a collision to reach their neighbors, which ultimately leads to the system reaching an absorbing state. Hence, after a collision, two particles will simply have a velocity $\Delta$ which on average is dissipated to zero over a distance $\Delta/\gamma$. This distance must be equal to the mean free path, which lets us derive the equation of the critical packing fraction:
@@
As expected, $\tau_{a}$ is proportional to $1/\gamma$. Our equations are now devoid of adjustable parameters. In the inset of Fig. 8a, we compare this new theory (Eqs. (18), (20) and (22)) with the simulations. The agreement is very satisfactory, especially in view of the assumption made (Gaussianity of the active particles, homogeneity, absence of spatial and temporal correlation, mean field approach, etc.). The temperatures in the inset are rescaled by $m\Delta^{2}/2$ to emphasize the fact that the system reaches an absorbing state when the energy of the active particles is close to the minimal excitation $m\Delta^{2}/2$. Indeed, the corrected theory predicts the transition when the naive theory –now representing the theoretical energy of the active particles only– is close to $m\Delta^{2}/2$. Three points however warrant attention. (i) The equations are solved in the mean field approximation, hence they cannot be exact close to the transition and this would be reflected, for example, in the discrepancy of critical exponents. (ii) With the new theory, the transition is not predicted to happen exactly at the critical packing fraction obtained from the mean free path argument $\phi_{c}^{MFP}$ even if it is used to find $\tau_{a}$. This is because the temperature of the active particles predicted Eq. (18) at $\phi_{c}^{MFP}$ is obviously not strictly equal to $m\Delta^{2}/2$. (iii) The coefficient of restitution $\alpha$ is close to 1, hence, the probability distribution is expected to be almost Gaussian in the active region where the drag is not playing a major role [134, 135]. The predictions would thus not be as good for systems with stronger dissipation at collision.
@@ IV.2 Synchronization dependent $\Delta$ model
Figure 9: a) Theoretical non-equilibrium potential $U$ as a function of $T$ for different $\phi$ and $\hat{\tau}/\tau_{s}=0.1$, $\hat{\tau}\Delta/\sigma=0.025$, $\hat{\tau}\gamma=0.01$ and $\alpha=0.95$. The potential behave exactly as an equilibrium one for a thermodynamic first order phase transition as the order parameter is changed. b) Effect of synchronization over the nature of transition and the critical packing fraction: $N=20000$, $\hat{\tau}\Delta/\sigma=0.025$, $\hat{\tau}\gamma=0.01$ and $\alpha=0.95$. Inset is a typical snapshot of the system close to the transition point.
@@
The model Eq. (7) implies that the value of $\Delta_{ij}$ depends on the time since the last collision for both particles involved $i$ and $j$. If at least one particle has travelled for a time larger than $\tau_{s}$, the value of $\Delta_{ij}$ is $\Delta$ otherwise it is set to 0. In this case an exact calculation of the collisional average in Eq. (11) is challenging, however we can make a reasonable estimation by assuming that the synchronization dependent $\Delta_{ij}$ can be replaced by an effective one: $\overline{\Delta}$. Here, we propose:
@@
As a qualitative check, when $01/\tau_{s}>$ 0, the non-equilibrium potential given by the integral of Eq. (25) has the same functional shape as an equilibrium Landau free energy [136, 137] of a system undergoing a first-order phase transition as seen in Fig. 9a. A quantitative comparison between the simulation and the theory is depicted in Fig. 9b. The theory, corresponding to the local minima of $U$ or equivalently the root of $G$ (found numerically), consistently predicts the observed transition. We note that the mean-field nature of the solution limits our ability to accurately determine the critical packing fraction. Near the transition point, density and spatial synchronization inhomogeneities arise due to the local nature of synchronization and due to the dissipative nature of the collisions between synchronized particles, as is depicted on the snapshot in the inset of Fig. 9b and in Fig. 7. We can see clusters of synchronized and asynchronized particles, which can reasonably explain the inaccuracies of the theories (especially close to the transition) together with the effective $\Delta$ approximation.
@@
where $K_{n/2}$ are positive numbers for $n\geq 6$. Like in previous investigations [52, 138, 139, 140, 141], the non-equilibrium potential depends on half-integer powers of $T$. This contrasts with the usual requirement of analyticity for equilibrium Landau free energies.
@@
playing a role akin to the $\epsilon$ of Ref. 31. When $K_{5/2}<0$, the dissipation exceeds the energy injection, leading to an easily unstable absorbing state with respect to small perturbations. Conversely, when $K_{5/2}>0$, the absorbing state is stable. We conclude that the critical packing fraction $\phi_{c}$ approximately satisfies $K_{5/2}(\phi_{c})=0$ when the drag can be neglected. Importantly, we see that for this transition, the critical packing fraction is approximately given by a competition between the average energy injected and dissipated at collision. Instead, for the continuous transition, the critical packing fraction was dictated by the competition between energy injection at collision and dissipation during the free flight. Notably, the mean free path argument (Eq. (21)) predicts a critical packing fraction for the continuous transition independent on $\alpha$. This is consistent with our understanding that synchronization allows the system to reach the absorbing state for a range of packing fractions that would have corresponded to the active phase in the asynchronized case. Increasing $\phi$, and in turn the collision rate $\omega(T,\phi)$, gives the grains less time to synchronize between collisions, which happens over a characteristic time $\tau_{s}$. For the coarse-grained model, we have numerically confirmed that the typical time between collisions $1/\omega$ is always on the same order as (but slightly below) $\tau_{s}$ in the first active state after the discontinuous transition.
This description in terms of potential also sheds light on the nucleation time analysis performed in Sec. III.1. The active state is never truly stable, even when the corresponding minimum is below the one of the absorbing state. At any packing fraction, a system that samples the absorbing state will never recover. Mathematically, this is understood through the multiplicative noise associated to the temperature [1, 142]. Indeed, to model fluctuations beyond the mean-field approximation (while still neglecting spatial dependence), we can introduce a stochastic term to Eq. (27), with the Itô prescription [143, 144, 145]:
@@
with $K$ a constant related to the intensity of the noise and $\eta$, a Gaussian white noise with 0 average and unit variance. The absorbing state is devoid of noise since $T_{a}=0$ and thus does not evolve.
Due to these peculiarities, it becomes questionable and potentially misleading to use equilibrium statistical physics terminology, such as metastable states, to describe states corresponding to a local minimum of the effective potential. For finite system sizes, the active state will invariably be metastable with respect to the absorbing state, irrespective of the value of $U$. Nonetheless, if there is no obvious special physics associated with the packing fraction at which the two minima are at equal height, transition time from the active state to the absorbing state can still be defined numerically as done in Sec. III.1 or theoretically through instantons and WKB approximations [146, 147, 148, 149]. These computations demonstrate that, even under the influence of multiplicative noise and out-of-equilibrium dynamics, a generalized Arrhenius law remains applicable in these systems. It provides an insight on what we call metastability within the context of absorbing phase transitions.

### v2

@@ IV Kinetic theory
In this section we develop the kinetic theory of the homogeneous state for the effective 2D model. We predict quantitatively the transition points and the steady state temperature of our systems. We relate our findings to results already obtained in the literature.
@@ IV.1 Simple $\Delta$ model
Figure 8: Comparison between the different theories applied to the simple $\Delta$ model without synchronization and simulations. The analysis is accompanied by examples of the breakdown of the assumptions used. a) Comparison of the theory of the simple model with the simulations. “Naive theo.” is Eq. (15). “Corrected theo.” is Eqs. (18) and (20). Dark dashed lines are the theoretical critical packing fractions derived from the mean free path argument. The inset is in a semi-log scale close to the transition and in units of $m\Delta^{2}/2$. We see that when the naive theory predicts a temperature of the order $m\Delta^{2}/2$, the system reaches an absorbing state in simulation. $N=5000$ and $\alpha=0.95$. The increasing values of $\Delta/\gamma\sigma$ are: 1.25, 1.5 and 1.75. b) Evolution of the velocity autocorrelation function as the critical density of the continuous transition is approached. The system starts to exhibit backscattering and the prediction from the Enskog theory with molecular chaos cannot be trusted. c) Evolution of the pre-colliding velocity $v_{coll}$ as a function of the proximity to the absorbing state. $v_{coll}$ tends to saturate, implying that the active particles have a finite energy at the transition. d) Evolution of the measured frequency of collision as the critical point is approached ($T^{ss}\rightarrow 0$) by varying $\Delta$ from $\Delta_{c}$, which gives a vanishing $T^{ss}$ to a large value where the measured frequency of collision matches the one predicted by the kinetic theory $\omega_{e}$. At small temperatures, the frequency of collision increases linearly with $T^{ss}$ because it becomes proportional to the number of active particles.
Before introducing the theory of the full effective 2D model, it is useful to discuss the scenario where particles never synchronize which corresponds to the case $\tau_{s}\rightarrow\infty$ and, consequently, $\Delta_{ij}=\Delta$ is constant, see Eq (7). If the theory is accurate it should predict, in this situation, a continuous transition. We first recall the theory developed in Ref. maire2024interplay and introduced in Ref. brito2013hydrodynamic.
@@
Using the collision rule of the simple $\Delta$ model (Eq. (5)) and averaging over collisions, the energy change per collision can be estimated explicitly [brito2013hydrodynamic]:
@@
where $m$ is the mass of a particle. This result is based on the assumption of uncorrelated velocities (molecular chaos) and Gaussianity of the velocity distribution. The detailed calculations required to arrive at Eq. (12) can be found in Ref. maire2024interplay. These assumptions immediately imply that the frequency of collision is given by the so-called Enskog frequency of collision $\omega_{e}$ [pagonabarraga2001randomly]:
@@
where $l(\phi)$ is the mean free path and $\chi$ is the Enskog factor taken to be the radial pair distribution function at contact of an equilibrium system at the same density. It arises from the Enskog’s correction to the molecular chaos assumption [pagonabarraga2001randomly]. It can be approximated for $\phi<0.6$ by [mulero2009equation]:
@@
In Fig. 8a, we compare this theoretical prediction with measurements done with simulations of the effective 2D model at fixed $\Delta$ (i.e. no synchronization) and indeed the theory seems to work very well. This result, however, must be taken with care since the simulations reach an absorbing state at a finite packing fraction and this behavior is not predicted by the theory, which predicts zero temperature only at zero packing fraction, see the inset of Fig. 8a. The assumptions of homogeneity, molecular chaos and Gaussian distribution of velocity do not hold close to or below the critical packing fraction due to the significant effect of viscous drag. As can be seen in Fig. 8b, close to the transition, molecular chaos is broken since the velocity autocorrelation function develops a negative minimum at a time roughly equal to the mean collision time that is a clear indication of a backscattering effect. This is incompatible with the assumption of uncorrelated or Markovian collisions underlying the molecular chaos assumption. Indeed, such a theory can only predict exponentially decaying velocity autocorrelation [van2001kinetic, gardiner2009stochastic, sarracino2010irreversible]. This backscattering is induced by the considerable slowdown of the particles before colliding with their neighbors and gaining velocities in a direction opposite to the one they came from. Treating this effect would require going beyond molecular chaos; an approach that is made possible, by some tedious calculation [van1998ring]. Moreover, we observed in Fig. 7 that close to the critical packing fraction, the dynamics is dominated by avalanches of active particles within a population of inactive particles. The transition could then occur via two different scenarios. Either the temperature of the active particles reaches 0 at $\phi_{c}$, or their number vanishes while they still possess a finite amount of energy. The second scenario is the one observed in our system, as illustrated in Fig. 8c. Indeed, as the critical point is approached, the pre-colliding velocity $v_{0}$ (the velocity just before a collision) tends to saturate. This indicated that, at the transition, the active particles have finite energy, roughly equal to the minimal excitation possible: $m\Delta^{2}/2$, but their number vanishes.
@@
The first term accounts for the conversion of inactive particles to active ones due to collisions. The second term is a rate of inactivation modulated by the probability for an active particle of not colliding in a time $\tau_{a}$ assuming that collisions are Poissonian [visco2008non]. It accounts for particles reaching the absorbing state due to an absence of collision between them and a neighbor in a time $\tau_{a}$. This $\tau_{a}$ is acting as a cutoff with a value that will be derived below and should be inversely proportional to the drag $\gamma$. We recall that $\omega/2$ is the number of collision events per particle happening in the system per unit of time, while $\omega$ is the average frequency at which one particle collides with others.
@@
with $W_{0}$ the principal branch of the Lambert $W$ function. $P_{a}^{ss}$ varies continuously, and the transition happens when the mean free time $1/\omega_{e}$ of the active particle is of the same order ($W_{0}(2)\simeq 0.85)$ as the stopping time $\tau_{a}$. It is gratifying to observe that Eq. (19) is the mean field equation describing the conserved directed percolation universality class, which is the correct equation to describe the vicinity of our phase transition [ma2023theory]. At this point, it is interesting to make a parallel with models on lattice. As said before, the criticality in our model comes from the vanishing of $P_{a}$ while the temperature of the active particles is fixed. This behavior closely resembles lattice models in the conserved directed percolation universality class, where the order parameter is the number of active particles and for which the activity kick allowing a particle to move on the lattice is fixed and finite [manna1990cascades, manna_sandpile_1999].
To close our equations, we must determine a suitable value for $\tau_{a}$. This can be achieved, for example, by finding an independent prediction for $\phi_{c}$. An argument presented in Ref. [lei2019hydrodynamics] allows us to find a lower bound for the critical packing fraction of the continuous transition. If we assume spatial homogeneity, at the critical packing fraction, particles must collide on average with zero velocity. Below this density, particles do not have sufficient momentum after a collision to reach their neighbors, which ultimately leads to the system reaching an absorbing state. Hence, after a collision, two particles will simply have a velocity $\Delta$ which on average is dissipated to zero over a distance $\Delta/\gamma$. This distance must be equal to the mean free path, which lets us derive the equation of the critical packing fraction:
@@
As expected, $\tau_{a}$ is proportional to $1/\gamma$. Our equations are now devoid of adjustable parameters. In the inset of Fig. 8a, we compare this new theory (Eqs. (18), (20) and (22)) with the simulations. The agreement is very satisfactory, especially in view of the assumption made (Gaussianity of the active particles, homogeneity, absence of spatial and temporal correlation, mean field approach, etc.). The temperatures in the inset are rescaled by $m\Delta^{2}/2$ to emphasize the fact that the system reaches an absorbing state when the energy of the active particles is close to the minimal excitation $m\Delta^{2}/2$. Indeed, the corrected theory predicts the transition when the naive theory –now representing the theoretical energy of the active particles only– is close to $m\Delta^{2}/2$. Three points however warrant attention. (i) The equations are solved in the mean field approximation, hence they cannot be exact close to the transition and this would be reflected, for example, in the discrepancy of critical exponents. (ii) With the new theory, the transition is not predicted to happen exactly at the critical packing fraction obtained from the mean free path argument $\phi_{c}^{MFP}$ even if it is used to find $\tau_{a}$. This is because the temperature of the active particles predicted Eq. (18) at $\phi_{c}^{MFP}$ is obviously not strictly equal to $m\Delta^{2}/2$. (iii) The coefficient of restitution $\alpha$ is close to 1, hence, the probability distribution is expected to be almost Gaussian in the active region where the drag is not playing a major role [garzo2018enskog, brey2013homogeneous]. The predictions would thus not be as good for systems with stronger dissipation at collision.
@@ IV.2 Synchronization dependent $\Delta$ model
Figure 9: a) Theoretical non-equilibrium potential $U$ as a function of $T$ for different $\phi$ and $1/(\gamma\tau_{s})=10$, $\Delta/(\gamma\sigma)=2.5$ and $\alpha=0.95$. The potential behave exactly as an equilibrium one for a thermodynamic first order phase transition as the order parameter is changed. b) Effect of synchronization over the nature of transition and the critical packing fraction: $N=20000$, $\Delta/(\gamma\sigma)=2.5$ and $\alpha=0.95$. Inset is a typical snapshot of the system close to the transition point.
@@
The model Eq. (7) implies that the value of $\Delta_{ij}$ depends on the time since the last collision for both particles involved $i$ and $j$. If at least one particle has traveled for a time smaller than $\tau_{s}$, the value of $\Delta_{ij}$ is $\Delta$ otherwise it is set to 0. In this case an exact calculation of the collisional average in Eq. (11) is challenging, however we can make a reasonable estimation by assuming that the synchronization dependent $\Delta_{ij}$ can be replaced by an effective one: $\overline{\Delta}$. Here, we propose:
@@
As a qualitative check, when $1/\tau_{s}>0$, the non-equilibrium potential given by the integral of Eq. (25) has the same functional shape as an equilibrium Landau free energy [goldenfeld2018lectures, cardy1996scaling] of a system undergoing a first-order phase transition as seen in Fig. 9a. A quantitative comparison between the simulation and the theory is depicted in Fig. 9b. The theory, corresponding to the local minima of $U$ or equivalently the root of $G$ (found numerically), consistently predicts the observed transition. We note that the mean-field nature of the solution limits our ability to accurately determine the critical packing fraction. Near the transition point, density and spatial synchronization inhomogeneities arise due to the local nature of synchronization and due to the dissipative nature of the collisions between synchronized particles, as is depicted on the snapshot in the inset of Fig. 9b and in Fig. 7. We can see clusters of synchronized and asynchronized particles, which can reasonably explain the inaccuracies of the theories (especially close to the transition) together with the effective $\Delta$ approximation.
@@
where $K_{n/2}$ are positive numbers for $n\geq 6$. [mari2022absorbing, PhysRevResearch.2.043390, 10.21468/SciPostPhys.8.5.074, PhysRevLett.130.207102, ptaszynski2025nonanalyticlandaufunctionalsshaping]
@@
playing a role akin to the $\epsilon$ of Ref. neel2014dynamics. When $K_{5/2}<0$, the dissipation exceeds the energy injection, leading to an easily unstable absorbing state with respect to small perturbations. Conversely, when $K_{5/2}>0$, the absorbing state is stable. We conclude that the critical packing fraction $\phi_{c}$ approximately satisfies $K_{5/2}(\phi_{c})=0$ when the drag can be neglected. Importantly, we see that for this transition, the critical packing fraction is approximately given by a competition between the average energy injected and dissipated at collision. Instead, for the continuous transition, the critical packing fraction was dictated by the competition between energy injection at collision and dissipation during the free flight. Notably, the mean free path argument (Eq. (21)) predicts a critical packing fraction for the continuous transition independent on $\alpha$. This is consistent with our understanding that synchronization allows the system to reach the absorbing state for a range of packing fractions that would have corresponded to the active phase in the asynchronized case. Increasing $\phi$, and in turn the collision rate $\omega(T,\phi)$, gives the grains less time to synchronize between collisions, which happens over a characteristic time $\tau_{s}$. For the effective 2D model, we have numerically confirmed that the typical time between collisions $1/\omega$ is always on the same order as (but slightly below) $\tau_{s}$ in the first active state after the discontinuous transition.
[hinrichsen2000non, dornic2005integration][van1992stochastic, canet2011general, honkonen2011ito]:
@@
[zakine2023minimum, elgart2004rare, meerson2011extinction, assaf2017wkb]